Barbara borrows $3500. She agrees to make monthly interest payments on the loan and will build up a sinking fund with monthly deposits to repay the principal with a single payment 21 months from now. If the interest being charged on the loan is j12 = 8% and the interest being earned on the sinking fund is j12= 5.2%, what is the monthly cost of the debt for Barbara?

I got 182.95 but is wrong. please help
Answer #1

Monthly interest paid on loan = amount borrowed * interest rate on loan / 12

Monthly interest paid on loan = $3,500 * 8% / 12 = $23.33

Monthly sinking fund payment is calculated using PMT function in Excel :

rate = 5.2%/12 (monthly rate earned on sinking fund = annual rate / 12)

nper = 21 (21 monthly deposits)

pv = 0 (beginning value of sinking fund is zero)

fv = -3500 (required value of sinking fund after 21 months)

PMT is calculated to be $159.56

X fac =PMT(5.2%/12,21,0,-3500) D E F B C A $159.56

Monthly cost of debt = Monthly interest paid on loan + Monthly sinking fund payment

Monthly cost of debt = $23.33 + $159.56

Monthly cost of debt = $182.89